Explain the procedures in the preparation of a 10% 25ml aqueous sodium chloride solution.
1
Expert's answer
2020-11-04T14:25:17-0500

Initially it's imrtant to understand the mass of sodium chloride to be used for this solution.

"Cm (NaCl)=\frac{m(solute)}{m(solute)+m(solvent)}⋅100%""%=frac\{};{}"

Assuming that 1 ml = 1 g (solution density is 1 g/ml)

"10=\frac{х}{25}⋅100"

х=2.5 g

Therefore, the mass of 2.5 g of NaCl has to mixed with 25 ml of water.